How much 60% antifreeze solution and 54% antifreeze solution should be combined to give 39 gallons of 58% antifreeze solution?
Answer by nerdybill(7384)

How much 60% antifreeze solution and 54% antifreeze solution should be combined to give 39 gallons of 58% antifreeze solution?
.
Let x = amount (gallons) of 60% solution
then
39-x = amount (gallons) of 54% solution
.
.60x + .54(39-x) = .58(39)
.60x + 21.06 - .54x = 22.62
.60x - .54x = 1.56
.06x = 1.56
x = 26 gallons of 60% solution
.
54% solution:
39-x = 39-26 = 13 gallons